OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = 2x2 − 3x + 0.5. What do the solutions of 0 = 2x2 − 3x + 0.5 represent on the graph of the equation? My Answer: Personally, I would use a graphing technology to find where the function f(x) = 2x^2 - 3x + 0.5 intersects the x-axis. The solutions of 0 = 2x^2 - 3x + 0.5 represent the x-intercepts of the graph. Does this sound okay?

OpenStudy (anonymous):

I have five more questions in this link that I wanted to get checked as well, is anyone is willing to take the time? http://answers.yahoo.com/question/index;_ylt=AmzVCpuDoLIuLXes5vZLu7jsy6IX;_ylv=3?qid=20120406181758AAkH1v9

OpenStudy (anonymous):

They are all word problems, so they may take time... depending on how good you are in math. I just want to make sure my answers are correct.

OpenStudy (anonymous):

sorry if it's to much

OpenStudy (anonymous):

Use complete sentences to explain how you would use graphing technology to find the solutions of 0 = 2x^2 − 3x + 0.5. I would plot the function and see where the graph hits the x axis . That points are soloutins

OpenStudy (anonymous):

oh, you answered it , and it is like my answer;

OpenStudy (anonymous):

yeah, I just need someone to check my answers, I already answered them :)

OpenStudy (anonymous):

I can do that

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

where are they?

OpenStudy (anonymous):

does that link for yahoo work?

OpenStudy (anonymous):

I posted them on yahooanswers, I posted the link

OpenStudy (anonymous):

first one is right

OpenStudy (anonymous):

yay!! :)

OpenStudy (anonymous):

second one should be 12 sec not 6sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that's funny, that was my original answer for that question, but then I guess I thought I did something wrong... thanks! :)

OpenStudy (anonymous):

3 is right

OpenStudy (anonymous):

These two I didn't post on yahoo only because I was pretty sure they were right, but while your here, no harm in checking right? :) 1. Alexei wants to hang a mirror in his boat but the mirror and frame must not have an area larger than 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The border must have equal width on each side.) 2x2 + 8x − 19.25 = 0 3x2 + 5x − 19.25 = 0 4x2 + 16x − 4.25 = 0<---- My Answer 4x2 + 16x + 15 = 0 2. Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach? 0.8 ft 1.6 ft 9.8 ft<---- My Answer 19.6 ft

OpenStudy (anonymous):

last one is 1.6 , not 0.8

OpenStudy (anonymous):

okay, i'll change that :)

OpenStudy (anonymous):

What about those last two? I'm pretty sure there correct?

OpenStudy (anonymous):

or at least I hope... lol :)

OpenStudy (anonymous):

last one is same as the one in yahoo

OpenStudy (anonymous):

there actually different, One is measuring ft and the other seconds

OpenStudy (anonymous):

I thought they were the same at first glance to but there actually not.. lol

OpenStudy (anonymous):

last question is correct, I got 9.8 feet as well

OpenStudy (anonymous):

awesome! :)

OpenStudy (anonymous):

other one is right as well

OpenStudy (anonymous):

yay!! thanks so much for your help!! :)

OpenStudy (anonymous):

I really do appreciate it!! :) I'm gonna log off no, I hope you have a good night! Thanks again! :)

OpenStudy (anonymous):

you too